The Hollings Corporation is considering a two-step buyout of the Norton Corporation. The latter firm has 2.5 million shares outstanding and its stock price is currently $40 per share. In the two-step buyout, Hollings will offer to buy 51 percent of Norton’s shares outstanding for $62 per share in cash and the balance in a second offer of 840,000 convertible preferred stock shares. Each share of preferred stock would be valued at 40 percent over the current value of Norton’s common stock. Mr. Green, a newcomer to the management team at Hollings, suggests that only one offer for all of Norton’s shares be made at $59.25 per share.

a. Calculate the total costs of the two alternatives. (Do not round intermediate calculations. Enter your answers in dollars, not millions (e.g., $123,456,000).)

Total Costs
Two step offer
Single offer

b. Which is better in terms of minimizing costs?

Two step offer
Single offer
Answer #1

a. Two step offer Total Costs = 51% shares cost + Preferred shares cost

= 2500000*51%*62 + 840000*40*(1+40%)

= $126090000

Single offer Total Costs = 2500000*59.25 = $148125000

b. Two step offer is better in terms of minimizing costs.